Calculate the pH after adding 10.00 mL of 0.15 M HCl to 50 mL of pure water.
Answer #1

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Given:

M1 = 0.15 M

V1 = 10 mL

V2 = 50 mL + 10 mL = 60 mL

use:

M1*V1 = M2*V2

M2 = (M1*V1)/V2

M2 = (0.15*10)/60

M2 = 0.025 M

This is concentration of HCl

So,

[H+] =0.025 M

use:

pH = -log [H+]

= -log (2.5*10^-2)

= 1.6021

Answer: 1.60